In a closed economy with no government, the consumption function is modeled by
C = 500 + 0.65Y. If investment falls by $200m, what is the change in national Y.
1
Expert's answer
2020-07-21T15:14:42-0400

If the consumption function is modeled by C = 500 + 0.65Y, then the money multiplier is:

"m = \frac{1} {1 - 0.65} = 2.86."

If investment falls by $200m, then the national income Y will decrease by "200\u00d72.86 =" $5.72m.